Working together, Larry, Moe, and Curly can paint an elephant in 3 minutes. Working alone, it would take Larry 10 minutes or Moe 6 minutes to paint the elephant. How long would it take Curly to paint the elephant if heworked alone?
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(30993)

Working together, Larry, Moe, and Curly can paint an elephant in 3 minutes. Working alone, it would take Larry 10 minutes or Moe 6 minutes to paint the elephant. How long would it take Curly to paint the elephant if heworked alone?
------------------
Larry paints 1/10 of the elephant per minute.
Moe paints 1/6 per minute.
Curly paints 1/x per minute, we don't know yet.
Together, they paint 1/3 per minute.
------------
1/10 + 1/6 + 1/x = 1/3
Multiply by 30
3 + 5 + 30/x = 10
30/x = 2
x = 15
It would take Curly 15 minutes.

Answer by ewatrrr(10682)


Hi
Let x represent curly's solo time in minutes
1/10 + 1/6 + 1/x = 1/3 |per minute being the equalizer
6x + 10x + 60 = 20x |mulitply each term by 60x so as all denominators = 1
60 = 4x
x = 15min, curly's solo time in minutes